The energy unit used by electrical utilities in th
Answer #1

1 Kwhr = 1 kw*1hr
= 1000 W * 3600 s
= 3.6*10^6 J

Answer: 3.6*10^6 J

1 mega joule = 1*10^6 J
= 1*10^6 J / (3.6*10^6 J/jwh)
= 0.278 Kwh

cost = 0.278 kwh * $ 0.09 per kwh
= $ 0.025

Answer: $ 0.025